OpenStudy (anonymous):

-3 + m/9=10 solve for m

OpenStudy (anonymous):

m = 117

OpenStudy (anonymous):

-3 turns to +3 and you add to both sides making it 13. Then you take 9 and instead of divide you multiply both sides by 9.?

OpenStudy (anonymous):

m/9=13 m=13*9

OpenStudy (anonymous):

thats not an option orphtroll

OpenStudy (anonymous):

A. -30 B. 63 C. 87 D. 93

ganeshie8 (ganeshie8):

\(\frac{-3+m}{9} = 10 \) -3 + m = 90 m = 93

OpenStudy (anonymous):

thankyou!